Find area of quadrilateral formed by lines 2|x|+3|y|=6

Grade:9

1 Answers

Yeesha Kalal
7 years ago
The 4 lines forming quadrilateral are 2x+3y=6, 2x-3y=6, 3y-2x=6, 2x+3y=(-6). On solving for their intersection points(by solving 2 eq. at a time) we get (0,2) (3,0) (-3,0) (0,-2) as vertices of the quadrilateral. If u make a diagram in co-ordinate system, by simple geometry, area is 12.